You own a wholesale plumbing supply store. The store currently generates revenues of $1.02 million per year. Next year, revenues will either decrease by 9.7% or increase by 4.7%, with equal probability and then stay at that level as long as you operate the store.

You own the store outright. Other costs run $890,000 per year. There are no costs of shutting down, in that case, you can always sell the store for $520,000. What is the business worth today if the cost of capital is fixed at 9.9%?
Answer #1

The value of the plumbing supply store is max of all future profits and 520,000 also the value of sales.

if there is a reduction in revenue it shod be shut down.

if revenue increases then the profits are this time at:

1087940-921060=166880 + 166880/0.1=1835680

Which is good to keep it live and running.